x-9y+4z=1
-2x+9y-4z=-3
2x+y-4z=-3
What does x, y, and z equal! I need to solve either by substitution or elimination

Found 2 solutions by TimothyLamb, stanbon:
Answer by TimothyLamb(4379)

x - 9y + 4z = 1
-2x + 9y - 4z = -3
2x + y - 4z = -3
---
put the system of linear equations into standard form
---
x - 9y + 4z = 1
-2x + 9y - 4z = -3
2x + y - 4z = -3
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 2
y = 1
z = 2
---
by elimination:
---
x - 9y + 4z = 1
-2x + 9y - 4z = -3
2x + y - 4z = -3
---
add equations 1 and 2:
---
x - 9y + 4z = 1
-2x + 9y - 4z = -3
---
x + 4z = 1
-2x - 4z = -3
---
-x = -2
x = 2
---
add equations 2 and 3:
---
-2x + 9y - 4z = -3
2x + y - 4z = -3
---
9y - 4z = -3
y - 4z = -3
---
10y - 8z = -6 (Equation A)
---
substitute x=2 into equation 1:
---
2 - 9y + 4z = 1
-9y + 4z = -1 (Equation B)
---
add equations A and 2*B
---
10y - 8z = -6
2(-9y + 4z = -1)
---
10y - 8z = -6
-18y + 8z = -2
---
-8y = -8
y = 1
---
substitute x=2 y=1 into equation 1:
---
2 - 9 + 4z = 1
4z = 8
z = 2
---
answer:
x = 2
y = 1
z = 2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php

Answer by stanbon(75887)

x-9y+4z=1
-2x+9y-4z=-3
2x+y-4z=-3
-------
Add 2 times the 1st row to the 2nd row
Subtract 2 times the 1st row from the 3rd row
-------------------------------------------------
x-9y+4z=1
0-9y+4z=-1
0+19y-12z=-5
---------------------
Add 3 times the 2nd row to the 3rd to get:
x-9y+4z=1
0-9y+4z=-1
0-8y=0=-8
----------------
Solve the 3rd equation for "y"::
Add 3 times the 2nd row to the 3rd to get:
x-9y+4z=1
0-9y+4z=-1
y = 1
------
Substitute y = 1 into the 2nd equation to solve for "z".
x-9y+4z=1
z = 2
y = 1
-----
Solve for "x":
x = 2
----------------
Cheers,
Stan H.
-----------------------